# CHOICE: Coordinated Human-Object Interaction in Cluttered Environments for Pick-and-Place Actions

### v1

@@ Abstract.
Animating human-scene interactions such as pick-and-place tasks in cluttered, complex layouts is a challenging task, with objects of a wide variation of geometries and articulation under scenarios with various obstacles. The main difficulty lies in the sparsity of the motion data compared to the wide variation of the objects and environments as well as the poor availability of transition motions between different tasks, increasing the complexity of the generalization to arbitrary conditions. To cope with this issue, we develop a system that tackles the interaction synthesis problem as a hierarchical goal-driven task. Firstly, we develop a bimanual scheduler that plans a set of keyframes for simultaneously controlling the two hands to efficiently achieve the pick-and-place task from an abstract goal signal such as the target object selected by the user. Next, we develop a neural implicit planner that generates guidance hand trajectories under diverse object shape/types and obstacle layouts. Finally, we propose a linear dynamic model for our DeepPhase controller that incorporates a Kalman filter to enable smooth transitions in the frequency domain, resulting in a more realistic and effective multi-objective control of the character. Our system can produce a wide range of natural pick-and-place movements with respect to the geometry of objects, the articulation of containers and the layout of the objects in the scene.
@@ 1. Introduction
Controlling characters to perform pick-and-place motions in cluttered environments with obstacles, such as furniture, has significant applications in video games, virtual reality (VR), and robotics. These environments, often found in household chores, can be surprisingly complex and challenging to automate compared to most paid day jobs. The intricacy of the environment escalates due to diverse object shapes, functionalities, and scene layouts with hierarchical levels of objects and varied spatial relations. Planning viable collision-free trajectories for interacting with household objects, like those found in a real-world kitchen with various surface types and tightly packed items, presents an immense challenge. Synthesizing motions for a full-body character that can effectively navigate and interact in such environments is a fundamental yet crucial issue that requires a well-structured approach.
@@
Existing methods that generate human-object interactions in an indoor scene operate in simplified environments. For example, when grasping and manipulating light weight objects, the target object is often located in an open space with no obstacles, where the character simply needs to step forward and grasp it  (Wu et al., 2022; Ghosh et al., 2023). In reality, however, a milk bottle could be contained in a closed refrigerator that is full of items, or a bowl may be contained in a drawer together with other dishes. A successful interaction with the target object become a multi-step process. First, the character needs to prepare an open environment, by opening the refrigerator or pulling out the drawer. Next, they need to clear a collision-free path for the hand to reach the target object, by moving or even taking out nearby obstacles. Finally, the target object may need to be taken out through a narrow opening without colliding with other objects or the environment. Throughout the entire process, the person will coordinate all parts of the body for stepping, opening and grasping tasks according to the layout of the obstacles: for example, a person might side-step and open the cabinet with one hand while reaching out for the target object with the other hand. Although the whole process involves a series of discrete subtasks, the motion is planned such that the entire process is efficient and transitions are smooth.
@@
To tackle these challenges, we propose an integrated system composed of three key modules: a neural implicit trajectory planner that plans collision-free wrist motions, a bimanual scheduler that provides a set of subgoals in terms of keyframes for achieving the goal, and a real-time character controller that synthesizes the motion to reach the planned trajectories and subgoals with smoothness in the DeepPhase (Starke et al., 2022) latent space. Our interaction system approaches the interaction synthesis problem as a hierarchical goal-driven task. Given the complexity of directly dealing with the abstract goal control signal, which only specifies the target object and action, we break down these abstract signals into three-key-joint goal sequences using the Bimanual Scheduler and Implicit Neural Trajectory Planner. We carefully design the inputs and outputs of each module to progressively transit the abstract and sparse goal towards temporally dense and full-body level motion.
@@
To construct a valid 6D trajectory of an object to be grasped and picked out/put in a cluttered environment with multiple obstacles, we develop a novel data-driven implicit neural representation from sparse motion capture dataset. Our approach is inspired by the recent implicit neural representations of 3D shapes (Park et al., 2019; Marschner et al., 2023), scenes (Mildenhall et al., 2020), and navigation (Li et al., 2021): Their success inspires us to develop a unified representation that describes the spatial-temporal relationship between objects in the environment and the grasping hand. The grasping and placing trajectories from our extensively captured motion dataset could be viewed as an an-isotropic level set propagation process. Using these data as training samples, and SDF of the objects and obstacles as the additional inputs, we construct an implicit time-of-arrival field whose gradient represents the trajectory of the hand to reach the target object at any point in the space. During inference, our implicit field robustly plans the trajectory functioned on unseen objects and furniture placements.
@@
As our motions involve a series of discrete subtasks by different parts of the body, a naive controller can result in unsmooth and unrealistic coordination and transitions when the interaction goal switches. Moreover, subtle and complex locomotion, such as side-stepping and body twisting, are coupled simultaneously with upper-body interactions. Reproducing such motion realism from demonstration is challenging. To cope with this issue, we build our interaction controller upon the DeepPhase model (Starke et al., 2022), which has demonstrated high motion quality for acyclic motions like dancing. It is well-known that the frequency domain signals can retain the motion details while simplifying the motion representation. We found that the phase provides a compact and efficient low-dimensional feature space to help formulate the full-body motion, serving as a suitable goal feature for control. We then simplify the control law, making it easier to learn after abstracting the motion into our designed DeepPhase state space. Furthermore, we propose a linear dynamic model for our DeepPhase  (Starke et al., 2022) controller that incorporates a Kalman filter to enable smooth state transitions in the frequency domain. This approach further bridges the sparse interaction instructions to continuous-time goals ahead, resulting in a more realistic and effective control of the character while always following the interaction constraints.
@@
To also plan the sequence of motions that are required for reaching the target object, we propose a scheduler for bi-manual coordination. By integrating all our proposed methods, we create a comprehensive interaction system.
@@
We demonstrate our system’s ability to synthesize realistic pick-and-place interactions in highly complex kitchen environments (see Fig. 1). The system can produce subtle movements such as stepping, turning, and manipulation co-temporarily done by two hands in quick response to the user instructions; the system can generalize to diverse scene layouts and shapes.
@@
  • A systematic method for synthesizing long-term interaction in complex environments, hierarchically integrating three innovative sub-modules with significant enhancements individually, verified through quantitative and qualitative experiments,
  • an extensive motion dataset that includes precise fully-body manipulation including the motions of fingers and multiple types of objects, in contextual lifelike indoor scenes, which consists of motions such as taking out/putting in objects from/into containers, manipulation such as pouring, taking off the cap, etc.,
  • a neural implicit trajectory planner that generates realistic and plausible grasping/placing hand trajectories according to the object type, location and layout of the environment,
  • a DeepPhase-based interaction controller that robustly tracks goals of two-hands and the hip and
  • a bimanual scheduler that produces a series of keyframes for the hip and the wrist for pick-and-place tasks of objects within/to be put in a various containers, including shelves, drawers and cabinets.
@@ 2. Related Works
The field of motion synthesis has recently seen rapid advancements, utilizing both data-driven and physics-based methodologies (Peng et al., 2021, 2022; Tessler et al., 2024). In the realm of human-scene interaction, kinematics-based humanoid controllers have proven effective in executing interactive motions such as sitting, opening doors, and carrying objects (Starke et al., 2019; Zhang et al., 2022). By incorporating generative frameworks such as conditional Variational Autoencoders (cVAE) (Hassan et al., 2021), Diffusion models (Pi et al., 2023; Cen et al., 2024), or by further learning a Reinforcement Learning (RL) policy to control the VAE’s learned latent space (Zhao et al., 2023a; Luo et al., 2023), a more diverse range of interaction behaviors can be synthesized. In the application of physics simulation, diverse interaction motions such as catching and carrying, coupled with locomotion skills, have been realized by distilling a variety of expert demonstrations and RL-based control with visual input (Merel et al., 2020). More recent progress involves incorporating adversarial training in imitation learning to synthesize realistic interactions (Hassan et al., 2023). While these approaches yield impressively realistic animations, they are often constrained by the type of objects involved or the complexity of the scene.
@@
Enhanced by the construction (Hassan et al., 2019) or capture (Araujo et al., 2023) of real-life scene and motion pairs, human interactions can be synthesized in more clustered and complex scenarios(Li et al., 2023a). By extending the PROX dataset (Hassan et al., 2019) with semantic labels and scene segmentation, interaction poses can be conveniently generated through semantic control (Zhao et al., 2022). To compensate for datasets that lack motion-scene pairs, LAMA (Lee and Joo, 2023) and DIMOS (Zhao et al., 2023b) are introduced to fulfill appropriate motions in the desired scene interaction by incorporating RL. Although these methods can adapt to clustered motions, most interactions are limited to static scenes, lacking details of dynamic objects and finger motions.
@@
To produce fine-grained finger motions for object interaction, a wide range of methods have been explored, including the introduction of canonicalized contact representations (Zheng et al., 2023), grasp fields (Karunratanakul et al., 2020), and hand-object spatial representations (Zhang et al., 2021). Leveraging a physics engine, contact sampling (Ye and Liu, 2012), and DRL-based control (Christen et al., 2022; Zhang et al., 2023; Hu et al., 2024) are utilized to synthesize physically plausible manipulations. While these methods focus on isolated hand-object interactions, recent approaches have learned to synthesize full-body motions driven by dynamic object movements (Li et al., 2023b, a). Incorporating part-wise motion priors (PMP) (Bae et al., 2023) also allows for the physical assembly of hand and body motions that adapt to complex object movements. More recently, high-quality full-body object manipulation motion datasets have been acquired (Taheri et al., 2020; Bhatnagar et al., 2022; Fan et al., 2023), providing opportunities to explore full-body grasping pose synthesis (Xu et al., 2023) and motion control (Taheri et al., 2022; Wu et al., 2022; Ghosh et al., 2023; Braun et al., 2024). While GOAL (Taheri et al., 2022), GRIP (Taheri et al., 2024), and SAGA (Wu et al., 2022) concentrate on synthesizing full-body approaching and grasping motions, human-object interactions are further enhanced by conditioning on human intention (Ghosh et al., 2023) or physics-based hierarchical control (Braun et al., 2024; Luo et al., 2024). However, existing full-body object interaction works are still limited to relatively short motion sequences and monotonous scenes. Drawing from our interaction dataset, we propose a method for synthesizing full-body motion that engages with a variety of scenarios, featuring an array of furniture and everyday objects. Furthermore, we demonstrate that our framework is capable of executing a series of complex tasks involving object and scene interactions, all under interactive user control.
@@
Our control framework design is inspired by a stream of research that has been proposed to discover the periodic motion feature. This includes the periodical feature describing the feet-ground contacting and controlling the character using the phase in the frequency domain (Holden et al., 2017; Zhang et al., 2018; Starke et al., 2021). DeepPhase (Starke et al., 2022) has further proposed the Periodical Auto-Encoder (PAE) that computes a low-dimensional representation of the spatial-temporal body motion. By running a non-linear transformation and FFT, PAE can encode the full-body joint-space motion into several channels of periodic latent vectors, forming a phase manifold in the frequency domain.
@@
The advantage of the approach is that it can spatio-temporarily align a wide range of cyclic and acyclic movements with different hand and leg movements to form a motion manifold in the frequency domain, and then allowing detailed auto-regressive control of the body through trajectories projected on the ground. Such a characteristic is attractive for our pick-and-place task that requires the combination of complex leg movements, e.g., side-stepping and turning, and co-temporal hand movements, e.g., opening the cupboard on one hand while grasping an object with the other hand, that also requires position-based control for guiding the end effectors to the reach the target while avoiding obstacles in the environment.
@@
Figure 2. Overview of our CHOICE System: Perceiving the involved objects around the clicking goal object with the action instruction from the keyboard, our system arrange tasks and motions to the empty hands, and match bi-manual goals based on the character state and the goal tasks. The matched hand goal priors are then re-planned by our trajectory planning sub-system to generate a trajectory of manipulation goals that fit the runtime environment. After planning a set of navigation goals together with the manipulation trajectory, our goal coordination arranges the goal of the key-joints and a goal phase prior to the current character. The second stage of our system is the goal-driven motion controller, it auto-regressively updates the character motion towards the goal of keyjoints, together with a loop of the Kalman filter which corrects and sets the goal phase features.
@@ 3. System Overview
  • hand-object trajectory planning, managed by the trajectory planning sub-system (brown) shown in  Fig. 2 and § 4.2;
  • bimanual interaction scheduling, managed by the goal matching sub-system (grey) and the goal coordination (cyan) shown in Fig. 2 and § 6.3; and
  • fully-body motion control, managed by the goal-driven control sub-system (blue) shown in Fig. 2 and § 5.
When the user clicks the object to pick or place, the path to walk towards the object is first planned, then the series of keyframes for the key joints (left/right wrist and hip) based on contacts are scheduled and finally, collision-free trajectories of the arms to conduct the motion is computed. The DeepPhase-based controller then synthesizes the motion to follow the scheduled sub-goals and planned trajectories in real-time in the phase manifold. The major mathematical variables involved in these three sub-systems are summarized in Table. 1
@@
4. Traj. Plan $\mathbf{x}$ Query position
@@
5.1 Features $\mathbf{T}_{t}^{key}$ 2s trajectory of a key joint pivot on time $t$
@@
5.2 Controller $\mathbf{X}_{t}^{p}$ Phase feature pivoted at $t$ as state
@@
6. Scheduler $\bm{p}$ Character’s current pose when matching

### v2

@@ Abstract.
Animating human-scene interactions such as picking and placing a wide range of objects with different geometries is a challenging task, especially in a cluttered environment where interactions with complex articulated containers are involved. The main difficulty lies in the sparsity of the motion data compared to the wide variation of the objects and environments, as well as the poor availability of transition motions between different actions, increasing the complexity of the generalization to arbitrary conditions. To cope with this issue, we develop a system that tackles the interaction synthesis problem as a hierarchical goal-driven task. Firstly, we develop a bimanual scheduler that plans a set of keyframes for simultaneously controlling the two hands to efficiently achieve the pick-and-place task from an abstract goal signal such as the target object selected by the user. Next, we develop a neural implicit planner that generates hand trajectories to guide reaching and leaving motions across diverse object shapes/types and obstacle layouts. Finally, we propose a linear dynamic model for our DeepPhase controller that incorporates a Kalman filter to enable smooth transitions in the frequency domain, resulting in a more realistic and effective multi-objective control of the character. Our system can synthesize a rich variety of natural pick-and-place movements that adapt to different object geometries, container articulations, and scene layouts.
@@ 1. Introduction
Controlling characters to perform pick-and-place motions in cluttered environments with obstacles, such as furniture, has significant applications in video games, virtual reality (VR), and robotics. These environments, often found in household chores, can be surprisingly complex and challenging to automate. The intricacy of the environment escalates due to diverse object shapes, functionalities, and scene layouts with hierarchical levels of objects and varied spatial relations. Planning viable collision-free trajectories for interacting with household objects, like those found in a real-world kitchen with various surface types and tightly packed items, presents an immense challenge. Synthesizing motions for a full-body character that can effectively navigate and interact in such environments is a fundamental yet crucial issue that requires a well-structured approach.
@@
Existing methods that generate human-object interactions in an indoor scene operate in simplified environments. For example, when grasping and manipulating lightweight objects, the target object is often located in an open space with no obstacles, where the character simply needs to step forward and grasp it  (Wu et al., 2022; Ghosh et al., 2023). In reality, however, a milk bottle could be contained in a closed refrigerator that is full of items, or a bowl may be contained in a drawer together with other dishes. A successful interaction with the target object becomes a multi-step process. First, the character needs to prepare an open environment by opening the refrigerator or pulling out the drawer. Next, they need to clear a collision-free path for the hand to reach the target object by moving or even taking out nearby obstacles. Finally, the target object may need to be taken out through a narrow opening without colliding with other objects or the environment. Throughout the entire process, the person will coordinate all parts of the body for stepping, opening, and grasping tasks according to the layout of the obstacles: For example, a person might side-step and open the cabinet with one hand while reaching out for the target object with the other hand. Although the whole process involves a series of discrete sub-tasks, the motion is planned such that the entire process is efficient and transitions are smooth.
@@
To tackle these challenges, we propose an integrated system composed of three key modules: a neural implicit trajectory planner that plans collision-free wrist motions, a bimanual scheduler that provides a set of subgoals in terms of keyframes for achieving the goal, and a real-time character controller that synthesizes the motion to reach the planned trajectories and subgoals with smoothness in the DeepPhase (Starke et al., 2022) latent space. Our interaction system approaches the interaction synthesis problem as a hierarchical goal-driven task. Given the complexity of directly dealing with the abstract goal control signals, which only specify the target object and action, we compose three-key-joint goal sequences based on these abstract signals using the bimanual scheduler and the implicit neural trajectory planner. We carefully design the inputs and outputs of each module to progressively transition the abstract and sparse goal towards temporally dense and full-body-level motion.
@@
To construct a valid 6D trajectory of an object to be grasped and picked out/put in a cluttered environment with multiple obstacles, we develop a novel data-driven implicit neural representation from sparse motion capture datasets. Our approach is inspired by the recent implicit neural representations of 3D shapes (Park et al., 2019; Marschner et al., 2023), scenes (Mildenhall et al., 2020), and navigation (Li et al., 2021). Their success inspires us to develop a unified representation that describes the spatial-temporal relationship between objects in the environment and the grasping hand. The grasping and placing trajectories from our extensively captured motion dataset could be viewed as an an-isotropic level set propagation process. Using these data as training samples, and SDF of the objects and obstacles as the additional inputs, we construct an implicit time-of-arrival field whose gradient represents the trajectory of the hand to reach the target object at any point in the space. During inference, our implicit field robustly plans the trajectory functioned by unseen objects and furniture placements.
@@
As our motions involve a series of discrete subtasks by different parts of the body, a naive controller can result in unsmooth and unrealistic coordination and transitions when the interaction goal switches. Moreover, subtle and complex locomotion, such as side-stepping and body twisting, is coupled simultaneously with upper-body interactions. Reproducing such motion realism from demonstrations is challenging. To cope with this issue, we build our interaction controller upon the DeepPhase model (Starke et al., 2022), which has demonstrated high motion quality for acyclic motions like dancing. It is well-known that the frequency domain signals can retain the motion details while simplifying the motion representation. We found that the phase provides a compact and efficient low-dimensional feature space to help formulate the full-body motion, serving as a suitable goal feature for control. We then simplify the control law, making it easier to learn after abstracting the motion into our designed DeepPhase state space. Furthermore, we propose a linear dynamic model for our DeepPhase controller that incorporates a Kalman filter to enable smooth state transitions in the frequency domain. This approach further bridges the sparse interaction instructions to continuous-time goals ahead, resulting in a more realistic and effective control of the character while always following the interaction constraints.
@@
To also plan the sequence of motions that are required for reaching the target object, we propose a scheduler for bimanual coordination. By integrating all our proposed methods, we create a comprehensive interaction system.
@@
We demonstrate our system’s ability to synthesize realistic pick-and-place interactions in highly complex kitchen environments (see Fig. 1). The system can produce subtle movements such as stepping, turning, and manipulation co-temporarily done by two hands in quick response to the user’s instructions. Also, it can generalize to diverse scene layouts and shapes.
@@
  • A systematic method for synthesizing long-term interaction in complex environments, hierarchically integrating three innovative sub-modules with significant enhancements individually, verified through quantitative and qualitative experiments;
  • an extensive motion dataset that includes precise full-body manipulation, including the motions of fingers and multiple types of objects, in contextual lifelike indoor scenes, which consists of motions such as taking out/putting in objects from/into containers, manipulation such as pouring, taking off the cap, etc.;
  • a neural implicit trajectory planner that generates realistic and plausible grasping/placing hand trajectories according to the object type and location, and the layout of the environment;
  • a DeepPhase-based interaction controller that robustly tracks the goals of two hands and the hip; and
  • a bimanual scheduler that produces a series of keyframes for the hip and the wrist for pick-and-place tasks of objects within/to be put in various containers, including shelves, drawers, and cabinets.
@@ 2. Related Works
The field of motion synthesis has recently seen rapid advancements, utilizing both data-driven and physics-based methodologies (Peng et al., 2021, 2022; Tessler et al., 2024). In the realm of human-scene interaction, kinematics-based humanoid controllers have proven effective in executing interactive motions such as sitting, opening doors, and carrying objects (Starke et al., 2019; Zhang et al., 2022). By incorporating generative frameworks such as conditional Variational Auto-encoders (cVAE) (Hassan et al., 2021), Diffusion models (Pi et al., 2023; Cen et al., 2024), or by further learning a Reinforcement Learning (RL) policy to control the VAE’s learned latent space (Zhao et al., 2023a; Luo et al., 2023), a more diverse range of interaction behaviors can be synthesized. In the application of physics simulation, diverse interaction motions such as catching and carrying, coupled with locomotion skills, have been realized by distilling a variety of expert demonstrations and RL-based control with visual input (Merel et al., 2020). More recent progress involves incorporating adversarial training in imitation learning to synthesize realistic interactions (Hassan et al., 2023). While these approaches yield impressively realistic animations, they are often constrained by the type of objects involved or the complexity of the scene.
@@
Enhanced by the construction (Hassan et al., 2019) or capture (Araujo et al., 2023) of real-life scenes and motion pairs, human interactions can be synthesized in more clustered and complex scenarios(Li et al., 2024). By extending the PROX dataset (Hassan et al., 2019) with semantic labels and scene segmentation, interaction poses can be conveniently generated through semantic control (Zhao et al., 2022). To compensate for datasets that lack motion-scene pairs, LAMA (Lee and Joo, 2023) and DIMOS (Zhao et al., 2023b) are introduced to fulfill appropriate motions in the desired scene interaction by incorporating RL. Although these methods can adapt to clustered motions, most interactions are limited to static scenes, lacking details of dynamic objects and finger motions.
@@
To produce fine-grained finger motions for object interaction, a wide range of methods have been explored, including the introduction of canonicalized contact representations (Zheng et al., 2023), grasp fields (Karunratanakul et al., 2020), and hand-object spatial representations (Zhang et al., 2021). Leveraging a physics engine, contact sampling (Ye and Liu, 2012), and DRL-based control (Christen et al., 2022; Zhang et al., 2024; Hu et al., 2024) are utilized to synthesize physically plausible manipulations. While these methods focus on isolated hand-object interactions, recent approaches have learned to synthesize full-body motions driven by dynamic object movements (Li et al., 2023, 2024). Incorporating part-wise motion priors (PMP) (Bae et al., 2023) also allows for the physical assembly of hand and body motions that adapt to complex object movements. More recently, high-quality full-body object manipulation motion datasets have been acquired (Taheri et al., 2020; Bhatnagar et al., 2022; Fan et al., 2023), providing opportunities to explore full-body grasping pose synthesis (Xu et al., 2023) and motion control (Taheri et al., 2022; Wu et al., 2022; Ghosh et al., 2023; Braun et al., 2024). While GOAL (Taheri et al., 2022), GRIP (Taheri et al., 2024), and SAGA (Wu et al., 2022) concentrate on synthesizing full-body approaching and grasping motions, human-object interactions are further enhanced by conditioning on human intention (Ghosh et al., 2023) or physics-based hierarchical control (Braun et al., 2024; Luo et al., 2024). However, existing full-body object interaction works are still limited to relatively short motion sequences and monotonous scenes. Drawing from our interaction dataset, we propose a method for synthesizing full-body motion that engages with a variety of scenarios, featuring an array of furniture and everyday objects. Furthermore, we demonstrate that our framework is capable of executing a series of complex tasks involving object and scene interactions, all under interactive user control.
@@
Our control framework design is inspired by a stream of research that has been proposed to discover the periodic motion feature. This includes the periodical feature describing the foot-ground contacting and controlling the character using the phase in the frequency domain (Holden et al., 2017; Zhang et al., 2018; Starke et al., 2021). DeepPhase (Starke et al., 2022) has further proposed the Periodical Auto-Encoder (PAE) that computes a low-dimensional representation of the spatial-temporal body motion. By running a non-linear transformation and FFT, PAE can encode the full-body joint-space motion into several channels of periodic latent vectors, forming a phase manifold in the frequency domain.
@@
The advantage of the approach is that it can spatio-temporally align a wide range of cyclic and acyclic movements with different hand and leg movements to form a motion manifold in the frequency domain, and then allows detailed auto-regressive control of the body through trajectories projected on the ground. Such a characteristic is attractive for our pick-and-place task that requires the combination of complex leg movements, e.g., side-stepping and turning, and co-temporal hand movements, e.g., opening the cupboard on one hand while grasping an object with the other hand, that also requires position-based control for guiding the end effectors to reach the target while avoiding obstacles in the environment.
@@
Figure 2. Overview of our CHOICE System: Perceiving the involved objects around the clicking goal object (pic.1) with the action instruction from the keyboard, our system arrange tasks and motions to the empty hands, and match bimanual goals based on the character state and the goal tasks (green block that outputs pic.2). The matched hand goal priors are then re-planned by our trajectory planning sub-system (in dashed brown) to generate a trajectory of manipulation goals that fit the runtime environment (pic.4). After planning a set of navigation goals (pic.3) alongside the manipulation trajectory, our goal coordination arranges the goal for the joints (pic.5) and sets a goal phase prior for the current character. The second major stage of our system is the goal-driven motion controller. It auto-regressively updates the character’s motion towards the goal of keyjoints, forming a loop with a Kalman filter, which corrects and sets the goal phase features.
@@ 3. System Overview
  • hand-object trajectory planning, managed by the trajectory planning sub-system (dashed brown in  Fig. 2), which is described in § 4;
  • full-body motion control, managed by the DeepPhase interaction control sub-system (dashed blue in Fig. 2), which is described in § 5; and
  • bimanual goal scheduling, managed together by the goal matching sub-system (dashed green in Fig. 2) and the goal coordination (grey in Fig. 2), which are described in § 6.
When the user clicks the object to pick or place, the path to walk towards the object is first planned, then the series of keyframes for the keyjoints (left/right wrist and hip) based on contacts are scheduled, and finally, collision-free trajectories of the arms to conduct the motion are computed. The DeepPhase-based controller then synthesizes the motion to follow the scheduled subgoals and planned trajectories in real-time in the phase manifold. The major mathematical variables involved in these three sub-systems are summarized in Tab. 1.
@@
4.
Traj. Plan
$\mathbf{x}$ Query position
@@
5.1
Features
$\mathbf{T}_{t}^{key}$ 2s trajectory of a key joint pivot on time $t$
@@
5.2
Controller
$\mathbf{X}_{t}^{p}$ Phase feature pivoted at $t$ as state
@@
6.
Scheduler
$\bm{p}$ Character’s current pose when matching